How can I prove that the sequence of irrational numbers is unbounded
1
Expert's answer
2020-03-13T10:35:35-0400

In general, the question is not true.

We have many counter examples.

Before giving counter example, recall some definition.

SEQUENCE: A sequence of real numbers is a function defined on the set "N=" {1,2,3,.......} of natural numbers whose range is contained in the set "R" of real number.

BOUNDED SEQUENCE: A sequence "X=(x_n)" of real number is said to be bounded if there exist a real number "M" ">0" such that "|x_n|\leqslant M" for all "n\in N" .

Example 1. Let "(x_n)=\sqrt{5}" is a sequence of irrational number bounded by "\sqrt{5}"

Example 2. Let "(x_n)=\sqrt{2}+\frac{1}{n}" is a sequence of irrational number bounded by "\sqrt{2}+1"